The function J1 defined by

J1(x) = ∑((((−1)^n).x^(2n+1))/(n! (n + 1)! 2)).2n+1

n=0


,


is called the Bessel function of order 1.

a) Find its domain.

b) Graph the first several partial sums on a common screen using programming software.


1
Expert's answer
2022-03-02T16:29:58-0500

a) the domain of the function is the interval of convergence of power series. Let "a_n=\frac{(-1)^{n}x^{2n+1}}{n!(n+1)!2^{2n+1}}" and find the "lim_{n\to\infin}|\frac{a_{n+1}}{a_n}|"

"|\frac{\frac{(-1)^{n+1}x^{2(n+1)+1}}{(n+1)!((n+1)+1)!2^{2(n+1)+1}}}{\frac{(-1)^{n}x^{2n+1}}{n!(n+1)!2^{2n+1}}}|=|\frac{x^{2n+3}}{(n+1)!(n+2)!2^{2n+3}}\times\frac{n!(n+1)!2^{2n+1}}{x^{2n+1}}|"

"=|\frac{x^2}{(n+1)n!(n+1)(n+2)!2^2}\times n!(n+1)!|"

"=|\frac{x^2}{4(n+1)(n+2)}|"

"\to0\ as\ n\ \to\infin"

By ratio test, Since "lim_{n\to\infin|\frac{a_{n+1}}{a_{n}}|}=0\le1\ \forall\ x" , The given series converges ,With this ,write write the interval which is exactly the domain of the function

"(-\infin,\infin)"

b)